Cells A1 through A10 contain the values shown in the table below. What does this formula evaluate to? 

= SUM(A1:A10) - MAXI A1:A5) 

3 1 4 6 2 0 5 7 6 2 

  • 0

  • 29 

  • 30 

  • 36 

  • nothing - there is a syntax error in one of the formulas
Answer #1

SUM(A1 : A10) means sum of all values of cells from 1 to 10.

= 3+1+4+6+2+0+5+7+6+2

= 36.

MAX(A1: A5) means maximum value among the cells from 1 to 5.

i.e 6 is maximum value.

Now,

SUM(A1:A10) - MAX(A1:A5) = 36 - 6 = 30

SO the correct answer is 30.